# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ V-B5 Other WF defense methods
Cherubin et al. [190] proposed two application-layer defenses: ALPaCA, a server-side defense that modifies webpage content to obscure traffic patterns, and LLaMA, a lightweight client-side defense that randomizes request orders and introduces delays. The proposed defenses significantly reduce the accuracy of WF attacks, demonstrating their potential effectiveness in enhancing privacy for users of anonymous web services. De et al. [113] proposed two lightweight defenses, TrafficSliver-Net and TrafficSliver-App, which leverage traffic splitting over multiple Tor entry nodes to limit the data observed by any single entry node and distort repeatable traffic patterns exploited by WF attacks. The proposed methods aim to infer the content of encrypted and anonymized connections by analyzing patterns in data flows. Jahani et al. [191] proposed a WF defense model by minimizing the autocorrelation property in network traffic instances, which is identified as a key factor enabling WF attacks.
@@
Henri et al. [192] proposed HyWF, which splits traffic between two direction networks using a random splitting strategy. For each website access, a random probability determines how packets are distributed between the networks. This randomness makes it difficult for an adversary to reconstruct the original traffic pattern. Later, Reuter et al. [193] presented Traffic Splitting for Tor, a defense against WF attacks that leverages multipath transmissions to obscure traffic patterns and thwart malicious entry nodes. The approach distributes Tor cells across multiple entry relays, disrupting identifiable timing and volume patterns while minimizing overhead.
@@
Zhang et al. [194] introduced Unsupervised and Adaptive Tor Website Fingerprinting (UAF), which is designed to address the challenge of domain shifts in Tor WF attacks. Traditional WF methods assume that training and testing data share similar distributions, which is unrealistic due to diverse user environments. UAF leverages multi-source domain adaptation (MDA) to align features between labeled source domains and unlabeled target domains without requiring target labels. It employs three trace representations (raw directions, raw timestamps, and directional timestamps) to retain discriminative information and combines source-specific classifiers based on trace length distributions.
@@
Ling et al. [195] proposed a genetic programming approach to search for effective dummy cell injection patterns that can mislead DL-based WF classifiers. The method involves randomly injecting dummy Tor cells into labeled traffic traces and using a fitness function to evaluate the effectiveness of the generated variant traces. A fitness function based on feature distance is designed to guide the search for successful variant traces. The function calculates the distance between the feature vector of a variant trace and the average feature vector of a target website. A mutation direction control mechanism is introduced to ensure efficient search by using a sliding window to monitor the progress of the search.
@@
The splitting of traffic in the heterogeneous network resources has emerged as a new defense mechanism to disrupt traffic characteristics. But the previous works are inefficient and costly. To address this challenge, Wang et al. [196] introduced PWR (Path Weighted Random), a multipath traffic splitting mechanism designed to counter WF attacks by leveraging heterogeneous network resources. PWR can dynamically distribute traffic across multiple paths based on real-time link conditions, such as delay, packet loss, and congestion, using a Dirichlet distribution to randomize path selection. This approach disrupts traffic patterns, making it harder for attackers to classify user behavior, while maintaining transmission efficiency through ordered packet scheduling and link redundancy padding.
@@
Gu et al. [197] presented an innovative and practical solution for defending against WF attacks, showcasing the potential of adversarial techniques in enhancing online privacy without incurring prohibitive costs or overhead. This research contributes significantly to the field of network security, specifically in protecting user privacy while browsing the web. Liu et al. [18] presented SMART, a lightweight defense that splits traffic across multiple Tor entry relays, disrupting statistical features used by WFAs. SMART offers a practical, efficient defense against WFAs by combining multi-path routing and redundancy. However, Tor enables anonymous communication but remains vulnerable to flow correlation attacks that match traffic patterns between its entry and exit points. MUFFLER [198] addresses these issues by dynamically reshaping egress traffic through real-time connection mapping without adding padding or delays. It effectively disrupts correlation attacks, achieving only 1% true positive rate at 1% false positive rate, with minimal bandwidth and lower latency—while maintaining compatibility with the existing Tor infrastructure.
@@ VI-A1 Webpage assumption
The webpage assumptions mainly refer to the types of webpages that users browse. These include the single-page assumption, static webpage, passive webpage, cache-disabled assumption, replicability, traffic parsing, and multi-tab assumption with a fixed overlap ratio [27]. These assumptions simplify the scenario, or overestimating the adversary’s capabilities gives the attacker an unrealistic advantage.
Single-page assumption. The single-page assumption refers to the common assumption by researchers that the observer monitors only a single page on a website, typically the homepage. Some researchers believe that monitoring just the homepage is sufficient to determine whether a user has visited the site. However, this assumption is considered unrealistic [35], as users are likely to browse both the homepage and internal pages of a website.
Static webpage. The static webpage assumption suggests avoiding WF attacks on sites where content changes frequently, such as news websites or video content platforms that update every few minutes [35]. Training a traffic classifier is often time-consuming, and when webpage content changes, the attacker must retrain the classifier using new browsing traffic, which is very costly. As a result, attacks typically target static webpages or websites with regularly updated content, such as government agency sites.
Passive webpage. The passive webpage assumption refers to the idea that when Tor users browse a webpage, all active content related to the page is either disabled or inactive. Aminuddin et al. [27] defined active content as traffic traces resulting from dynamically loaded elements such as Flash, Java Applets, or JavaScript. Additionally, ads, plugins, and special modifications to the homepage during holidays should also be disabled, as such dynamic content can cause the traffic pattern of the same webpage to differ with each load.
@@ VI-A2 Concept drift
Juarez et al. [6] are the first to reveal the impact of website content updates on WF attacks. They observed that attack accuracy dropped by 50% in less than 10 days and fell to nearly zero after 90 days. Attarian et al. [199] conducted a systematic survey of data stream mining algorithms for handling concept drift, comparing approaches such as Adaptive Hoeffding Tree [199], Concept-Adapting Very Fast Decision Tree (CVFDT) [200], OzaBag [201], and Extreme Fast Decision Tree (EFDT) [202]. Among these, the Adaptive Hoeffding Tree achieved the highest accuracy by automatically detecting drift and updating its parameters. Building on this, Attarian et al. [64] proposed the Adaptive Online WF Attack (AdaWFPA), which models WF using the Adaptive Hoeffding Tree and incrementally updates the classifier with each new training instance. This design enables continuous adaptation to evolving website versions and sustained performance over time. However, since the model is updated continuously, it cannot be directly compared with the drift resilience of existing deep learning–based approaches.
Wang et al. [124] introduced snapshot WF (snWF), an ensemble method leveraging neural networks. Their snapshot ensemble strategy employs multiple snapshot models during testing, with final predictions obtained by averaging across models. The study further showed that concept drift has an even stronger impact in open-world scenarios, where performance degradation is more severe. Thus, evaluating drift in realistic open-world environments remains an important research challenge.
@@ VI-A3 User behavior assumption
The closed-world assumption dictates that a Tor user will only visit websites from a predetermined, finite set that an attacker has already collected. Researchers use this simplified scenario to train and test classifiers, which often leads to remarkably high accuracy rates. While useful for comparing the performance of different models, this controlled environment doesn’t reflect the open-ended nature of the real internet, where users can browse an infinite number of sites. Consequently, a classifier trained in a closed world is likely to perform poorly in a real-world setting where it encounters traffic from unknown websites [27].
The sequential browsing assumption simplifies user behavior by assuming that a user loads one webpage at a time and only moves to the next page after the previous one has fully loaded [37]. This eliminates any overlap in traffic from different websites. In reality, due to Tor’s slower connection speeds, users commonly open multiple browser tabs simultaneously. This concurrent browsing behavior generates a chaotic mix of overlapping traffic from different sites, making it incredibly difficult to isolate and identify the traffic of a single webpage.
@@ VI-B1 Dataset limitation
Collecting fresh, diverse, and reliable datasets is a major challenge for website fingerprinting. Attackers need up-to-date data to maintain high accuracy, but this is a resource-intensive process. Consequently, resource-constrained attackers must compromise by either limiting the number of monitored websites, using outdated data, or reducing the number of training samples per site [26]. Two primary methods for dataset collection have emerged, each with its own trade-offs: the snowball method and the gold panning method [26].
@@ VI-B2 WF attack model building
The fundamental goal of a classifier is to achieve high effectiveness, which can be measured using metrics such as accuracy, precision, and recall. Most existing studies evaluate classifier effectiveness in the closed-world setting [203], where traffic is assumed to originate from a limited number of targeted sources (e.g., specific websites). However, a classifier that performs well in a closed-world scenario often loses effectiveness in the more realistic open-world setting, where encrypted traffic originates from a much larger number of untargeted sources. In this context, targeted traffic may represent only a small fraction of the total traffic, and the classifier must be capable of reliably distinguishing targeted traffic from untargeted traffic. Furthermore, existing defense methods negatively impact classifier effectiveness because fewer discriminative features can be extracted from protected traffic. A key challenge, therefore, is to investigate how to maintain classifier effectiveness in distinguishing between targeted and untargeted traffic when all traffic is subject to protection mechanisms.
@@
In Section V, we sequentially presented the defenses against prevailing attacks as well as those specifically targeting Tor. A clear observation is that defenses in anonymity networks often lag behind emerging attacks. This phenomenon likely arises because defenders are frequently unaware of what threats to anticipate in advance. In Section V-B, we reviewed multiple defenses [35, 37, 7] against WF attacks that have been introduced successively over time. We also observed that anonymity analyses are often conducted under specific, constrained conditions. Comprehensive conclusions, therefore, can typically only be drawn when a wider range of real-world factors is incorporated into the analysis methods. This situation serves as a reminder for network designers and operators to exercise caution and employ rigorous methodologies in formal security assessments. To provide more robust defenses for anonymity networks, it is thus necessary to address the following aspects.
@@ VI-C1 Trade-Offs Under Multiple Factors
Technically, defense schemes proposed for anonymity networks often involve trade-offs. WF defenses, for example, typically need to balance fingerprinting accuracy, bandwidth overhead, and communication latency. This presents a challenge for designers and operators striving to build scalable, low-latency anonymity networks. Similarly, designing an ideal path selection algorithm that achieves load balancing, defends against both passive and active state-level adversaries, and minimizes information leakage is difficult to realize. For instance, Sun et al. [204] proposed a solution to defend against BGP hijacking and interception attacks. Later, Hanley et al. [205] improved this scheme by considering worst-case information leakage. However, Mitseva et al. [206] raised concerns about the effectiveness of the solution in [205] for defending against BGP attacks and reducing information leakage. In such scenarios, adaptive and tunable defense schemes may offer a practical alternative.
@@ VI-C2 Unreliable Evaluations of Defense Schemes
We have also observed the problem of unreliable evaluations of defense schemes within anonymity networks, particularly regarding weaknesses in WF defenses. Notably, many WF defenses overlook adversarial training, which is a critical factor. If a defense scheme is deterministic, an adversary with relevant knowledge could potentially reverse-engineer it. Furthermore, many WF defenses lack comprehensive evaluation metrics and often ignore important factors such as information leakage [207].
@@ VI-D3 Enhancing Data Diversity
Additionally, existing work lacks a systematic study of web caching strategies and often simplifies assumptions by disabling caching altogether. Karunanayake et al. [208] found that, under the influence of Tor browser caching, reloading a webpage produces traffic patterns that differ significantly from those generated during the first visit. This finding challenges the effectiveness of current attack models. Thus, future data collection efforts should include webpages with diverse caching strategies, allowing for deeper analysis of how caching behavior affects the performance of WF attacks.
@@ VI-D6 Large Language Models (LLMs) in WF attacks and defenses
LLMs are advanced deep learning models trained on massive text corpora to understand and generate natural language. Built on transformer architectures, LLMs capture complex contextual and sequential dependencies, enabling them to generalize across diverse data modalities beyond text. Recently, researchers have explored applying LLMs to WF [209, 210], where encrypted traffic traces are represented as symbolic sequences. In the LLM-based paradigm, encrypted traffic traces are converted into symbolic or textual sequences that capture packet-level or burst-level characteristics. LLMs can be integrated into WF tasks in multiple ways:

### v2

@@ V-B5 Other WF defense methods
Cherubin et al. [cherubin2017website] proposed two application-layer defenses: ALPaCA, a server-side defense that modifies webpage content to obscure traffic patterns, and LLaMA, a lightweight client-side defense that randomizes request orders and introduces delays. The proposed defenses significantly reduce the accuracy of WF attacks, demonstrating their potential effectiveness in enhancing privacy for users of anonymous web services. De et al. [de2020trafficsliver] proposed two lightweight defenses, TrafficSliver-Net and TrafficSliver-App, which leverage traffic splitting over multiple Tor entry nodes to limit the data observed by any single entry node and distort repeatable traffic patterns exploited by WF attacks. The proposed methods aim to infer the content of encrypted and anonymized connections by analyzing patterns in data flows. Jahani et al. [jahani2020effective] proposed a WF defense model by minimizing the autocorrelation property in network traffic instances, which is identified as a key factor enabling WF attacks.
@@
Henri et al. [henri2020protecting] proposed HyWF, which splits traffic between two direction networks using a random splitting strategy. For each website access, a random probability determines how packets are distributed between the networks. This randomness makes it difficult for an adversary to reconstruct the original traffic pattern. Later, Reuter et al. [reuter2021traffic] presented Traffic Splitting for Tor, a defense against WF attacks that leverages multipath transmissions to obscure traffic patterns and thwart malicious entry nodes. The approach distributes Tor cells across multiple entry relays, disrupting identifiable timing and volume patterns while minimizing overhead.
@@
Zhang et al. [zhang2021rap] introduced Unsupervised and Adaptive Tor Website Fingerprinting (UAF), which is designed to address the challenge of domain shifts in Tor WF attacks. Traditional WF methods assume that training and testing data share similar distributions, which is unrealistic due to diverse user environments. UAF leverages multi-source domain adaptation (MDA) to align features between labeled source domains and unlabeled target domains without requiring target labels. It employs three trace representations (raw directions, raw timestamps, and directional timestamps) to retain discriminative information and combines source-specific classifiers based on trace length distributions.
@@
Ling et al. [ling2022towards] proposed a genetic programming approach to search for effective dummy cell injection patterns that can mislead DL-based WF classifiers. The method involves randomly injecting dummy Tor cells into labeled traffic traces and using a fitness function to evaluate the effectiveness of the generated variant traces. A fitness function based on feature distance is designed to guide the search for successful variant traces. The function calculates the distance between the feature vector of a variant trace and the average feature vector of a target website. A mutation direction control mechanism is introduced to ensure efficient search by using a sliding window to monitor the progress of the search.
@@
The splitting of traffic in the heterogeneous network resources has emerged as a new defense mechanism to disrupt traffic characteristics. But the previous works are inefficient and costly. To address this challenge, Wang et al. [wang2023pwr] introduced PWR (Path Weighted Random), a multipath traffic splitting mechanism designed to counter WF attacks by leveraging heterogeneous network resources. PWR can dynamically distribute traffic across multiple paths based on real-time link conditions, such as delay, packet loss, and congestion, using a Dirichlet distribution to randomize path selection. This approach disrupts traffic patterns, making it harder for attackers to classify user behavior, while maintaining transmission efficiency through ordered packet scheduling and link redundancy padding.
@@
Gu et al. [gu2023online] presented an innovative and practical solution for defending against WF attacks, showcasing the potential of adversarial techniques in enhancing online privacy without incurring prohibitive costs or overhead. This research contributes significantly to the field of network security, specifically in protecting user privacy while browsing the web. Liu et al. [liu2023smart] presented SMART, a lightweight defense that splits traffic across multiple Tor entry relays, disrupting statistical features used by WFAs. SMART offers a practical, efficient defense against WFAs by combining multi-path routing and redundancy. However, Tor enables anonymous communication but remains vulnerable to flow correlation attacks that match traffic patterns between its entry and exit points. MUFFLER [seo2025muffler] addresses these issues by dynamically reshaping egress traffic through real-time connection mapping without adding padding or delays. It effectively disrupts correlation attacks, achieving only 1% true positive rate at 1% false positive rate, with minimal bandwidth and lower latency—while maintaining compatibility with the existing Tor infrastructure.
@@ VI-A1 Webpage assumption
The webpage assumptions mainly refer to the types of webpages that users browse. These include the single-page assumption, static webpage, passive webpage, cache-disabled assumption, replicability, traffic parsing, and multi-tab assumption with a fixed overlap ratio [aminuddin2023rise]. These assumptions simplify the scenario, or overestimating the adversary’s capabilities gives the attacker an unrealistic advantage.
Single-page assumption. The single-page assumption refers to the common assumption by researchers that the observer monitors only a single page on a website, typically the homepage. Some researchers believe that monitoring just the homepage is sufficient to determine whether a user has visited the site. However, this assumption is considered unrealistic [wang2013improved], as users are likely to browse both the homepage and internal pages of a website.
Static webpage. The static webpage assumption suggests avoiding WF attacks on sites where content changes frequently, such as news websites or video content platforms that update every few minutes [wang2013improved]. Training a traffic classifier is often time-consuming, and when webpage content changes, the attacker must retrain the classifier using new browsing traffic, which is very costly. As a result, attacks typically target static webpages or websites with regularly updated content, such as government agency sites.
Passive webpage. The passive webpage assumption refers to the idea that when Tor users browse a webpage, all active content related to the page is either disabled or inactive. Aminuddin et al. [aminuddin2023rise] defined active content as traffic traces resulting from dynamically loaded elements such as Flash, Java Applets, or JavaScript. Additionally, ads, plugins, and special modifications to the homepage during holidays should also be disabled, as such dynamic content can cause the traffic pattern of the same webpage to differ with each load.
@@ VI-A2 Concept drift
Juarez et al. [juarez2014critical] are the first to reveal the impact of website content updates on WF attacks. They observed that attack accuracy dropped by 50% in less than 10 days and fell to nearly zero after 90 days. Attarian et al. [pfahringer2007new] conducted a systematic survey of data stream mining algorithms for handling concept drift, comparing approaches such as Adaptive Hoeffding Tree [pfahringer2007new], Concept-Adapting Very Fast Decision Tree (CVFDT) [hulten2001mining], OzaBag [oza2001online], and Extreme Fast Decision Tree (EFDT) [manapragada2018extremely]. Among these, the Adaptive Hoeffding Tree achieved the highest accuracy by automatically detecting drift and updating its parameters. Building on this, Attarian et al. [attarian2019adawfpa] proposed the Adaptive Online WF Attack (AdaWFPA), which models WF using the Adaptive Hoeffding Tree and incrementally updates the classifier with each new training instance. This design enables continuous adaptation to evolving website versions and sustained performance over time. However, since the model is updated continuously, it cannot be directly compared with the drift resilience of existing deep learning–based approaches.
Wang et al. [wang2022snwf] introduced snapshot WF (snWF), an ensemble method leveraging neural networks. Their snapshot ensemble strategy employs multiple snapshot models during testing, with final predictions obtained by averaging across models. The study further showed that concept drift has an even stronger impact in open-world scenarios, where performance degradation is more severe. Thus, evaluating drift in realistic open-world environments remains an important research challenge.
@@ VI-A3 User behavior assumption
The closed-world assumption dictates that a Tor user will only visit websites from a predetermined, finite set that an attacker has already collected. Researchers use this simplified scenario to train and test classifiers, which often leads to remarkably high accuracy rates. While useful for comparing the performance of different models, this controlled environment doesn’t reflect the open-ended nature of the real internet, where users can browse an infinite number of sites. Consequently, a classifier trained in a closed world is likely to perform poorly in a real-world setting where it encounters traffic from unknown websites [aminuddin2023rise].
The sequential browsing assumption simplifies user behavior by assuming that a user loads one webpage at a time and only moves to the next page after the previous one has fully loaded [wang2016realistically]. This eliminates any overlap in traffic from different websites. In reality, due to Tor’s slower connection speeds, users commonly open multiple browser tabs simultaneously. This concurrent browsing behavior generates a chaotic mix of overlapping traffic from different sites, making it incredibly difficult to isolate and identify the traffic of a single webpage.
@@ VI-B1 Dataset limitation
Collecting fresh, diverse, and reliable datasets is a major challenge for website fingerprinting. Attackers need up-to-date data to maintain high accuracy, but this is a resource-intensive process. Consequently, resource-constrained attackers must compromise by either limiting the number of monitored websites, using outdated data, or reducing the number of training samples per site [shen2022machine]. Two primary methods for dataset collection have emerged, each with its own trade-offs: the snowball method and the gold panning method [shen2022machine].
@@ VI-B2 WF attack model building
The fundamental goal of a classifier is to achieve high effectiveness, which can be measured using metrics such as accuracy, precision, and recall. Most existing studies evaluate classifier effectiveness in the closed-world setting [shen2019webpage], where traffic is assumed to originate from a limited number of targeted sources (e.g., specific websites). However, a classifier that performs well in a closed-world scenario often loses effectiveness in the more realistic open-world setting, where encrypted traffic originates from a much larger number of untargeted sources. In this context, targeted traffic may represent only a small fraction of the total traffic, and the classifier must be capable of reliably distinguishing targeted traffic from untargeted traffic. Furthermore, existing defense methods negatively impact classifier effectiveness because fewer discriminative features can be extracted from protected traffic. A key challenge, therefore, is to investigate how to maintain classifier effectiveness in distinguishing between targeted and untargeted traffic when all traffic is subject to protection mechanisms.
@@
In Section V, we sequentially presented the defenses against prevailing attacks as well as those specifically targeting Tor. A clear observation is that defenses in anonymity networks often lag behind emerging attacks. This phenomenon likely arises because defenders are frequently unaware of what threats to anticipate in advance. In Section V-B, we reviewed multiple defenses [wang2013improved, wang2016realistically, cai2012touching] against WF attacks that have been introduced successively over time. We also observed that anonymity analyses are often conducted under specific, constrained conditions. Comprehensive conclusions, therefore, can typically only be drawn when a wider range of real-world factors is incorporated into the analysis methods. This situation serves as a reminder for network designers and operators to exercise caution and employ rigorous methodologies in formal security assessments. To provide more robust defenses for anonymity networks, it is thus necessary to address the following aspects.
@@ VI-C1 Trade-Offs Under Multiple Factors
Technically, defense schemes proposed for anonymity networks often involve trade-offs. WF defenses, for example, typically need to balance fingerprinting accuracy, bandwidth overhead, and communication latency. This presents a challenge for designers and operators striving to build scalable, low-latency anonymity networks. Similarly, designing an ideal path selection algorithm that achieves load balancing, defends against both passive and active state-level adversaries, and minimizes information leakage is difficult to realize. For instance, Sun et al. [sun2017counter] proposed a solution to defend against BGP hijacking and interception attacks. Later, Hanley et al. [hanley2019dpselect] improved this scheme by considering worst-case information leakage. However, Mitseva et al. [mitseva2023security] raised concerns about the effectiveness of the solution in [hanley2019dpselect] for defending against BGP attacks and reducing information leakage. In such scenarios, adaptive and tunable defense schemes may offer a practical alternative.
@@ VI-C2 Unreliable Evaluations of Defense Schemes
We have also observed the problem of unreliable evaluations of defense schemes within anonymity networks, particularly regarding weaknesses in WF defenses. Notably, many WF defenses overlook adversarial training, which is a critical factor. If a defense scheme is deterministic, an adversary with relevant knowledge could potentially reverse-engineer it. Furthermore, many WF defenses lack comprehensive evaluation metrics and often ignore important factors such as information leakage [li2018measuring].
@@ VI-D3 Enhancing Data Diversity
Additionally, existing work lacks a systematic study of web caching strategies and often simplifies assumptions by disabling caching altogether. Karunanayake et al. [karunanayake2023exploring] found that, under the influence of Tor browser caching, reloading a webpage produces traffic patterns that differ significantly from those generated during the first visit. This finding challenges the effectiveness of current attack models. Thus, future data collection efforts should include webpages with diverse caching strategies, allowing for deeper analysis of how caching behavior affects the performance of WF attacks.
@@ VI-D6 Large Language Models (LLMs) in WF attacks and defenses
LLMs are advanced deep learning models trained on massive text corpora to understand and generate natural language. Built on transformer architectures, LLMs capture complex contextual and sequential dependencies, enabling them to generalize across diverse data modalities beyond text. Recently, researchers have explored applying LLMs to WF [jiao2025llm, song2025redefining], where encrypted traffic traces are represented as symbolic sequences. In the LLM-based paradigm, encrypted traffic traces are converted into symbolic or textual sequences that capture packet-level or burst-level characteristics. LLMs can be integrated into WF tasks in multiple ways: